# The coarse Baum-Connes conjecture with filtered coefficients and product metric spaces

### v1

@@ 4. Product metric spaces
In this section, we will prove the coarse Baum-Connes conjecture with filtered coefficients for the product metric space of two metric spaces providing each of them satisfies the conjecture with filtered coefficients.
In this subsection, we build a bridge called the localization algebra along one direction with filtered coefficients connecting the coarse Baum-Connes conjecture with filtered coefficients for products and the conjecture with filtered coefficients for one component.
@@ 4.1. The coarse Baum-Connes conjecture along one direction with filtered coefficients
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two proper metric spaces with two countable dense subsets $Z_{X}$ and $Z_{Y}$, respectively. Let $T$ be an operator acting on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H$. The propagation along $X$ of $T$, denoted by $\operatorname{prop}_{X}(T)$, is defined to be
@@
Let $X$ and $Y$ be as above, the localization algebra along $X$ with filtered coefficients in $A$ of $X\times Y$, denoted by $C^{\ast}_{L,X,f}(X\times Y,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(X\times Y,A)$ satisfying that
@@
Similar definitions as above are appeared in [1][3].
Let $X$, $X^{\prime}$ and $Y$ be three proper metric spaces. For a uniformly continuous coarse map $g:X\rightarrow X^{\prime}$, then $g\times Id:X\times Y\rightarrow X^{\prime}\times Y$ is also a uniformly continuous coarse map. Thus by Lemma 3.13, there exists a $\ast$-homomorphism $Ad_{g\times Id}:C^{\ast}_{L,f}(X\times Y,A)\rightarrow C^{\ast}_{L,f}(X^{\prime}\times Y,A)$ defined by $V_{g\times Id}$. Moreover, for $u\in C^{\ast}_{L,X,f}(X\times Y,A)$, we have that
@@
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2},X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2},X_{1}\cap X_{2}$, respectively, then there is a Mayer-Vietoris six-term exact sequence
${K_{0}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{0}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(C^{\ast}_{L,{X},f}((X)\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,{X},f}((X)\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{1}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A)),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
Let $X$, $Y$ be two proper metric spaces and $N_{X}$ be a locally finite net in $X$. Then the evaluation at zero map $e$ induces the following isomorphisms
@@
Let $g:Y\rightarrow Y^{\prime}$ be a coarsely equivalent map, then $Id\times g:P_{k}(N_{X})\times Y\rightarrow P_{k}(N_{X})\times Y^{\prime}$ is also coarsely equivalent map which is uniformly continuous on $P_{k}(N_{X})$. Then by Corollary 3.8, we can show that $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y,A))$ are naturally isomorphic to $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y^{\prime},A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y^{\prime},A))$, respectively, for any $k\in\mathbb{N}$. Thus we completed the proof. ∎
Recall from Lemma 3.9 that $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration induced by the propagation on $Y$ and the filtration on $A$. We have the following relation between the coarse Baum-Connes conjecture along $X$ with filtered coefficients and the coarse Baum-Connes conjecture with filtered coefficients.
@@
Let $H$ be a separable Hilbert space and $U:H\rightarrow H\otimes H$ be a unitary. Fixed a unit vector $h_{0}\in H$ and let $P:H\rightarrow H,h\mapsto\langle h,h_{0}\rangle h_{0}$ which is a rank-one projection. The algebra $C_{f}^{\ast}(X\times Y,A)$ acts on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ and the algebra $C^{\ast}_{f}(X,C^{\ast}(Y,A))$ acts on $\ell^{2}(Z_{X})\otimes H\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$, where $Z_{X}$ and $Z_{Y}$ are two countable dense subsets in $X$ and $Y$, respectively. Define two $\ast$-homomorphisms:
@@
where $V:H\rightarrow H\otimes H$ maps $h$ to $h\otimes h_{0}$. Since the operator $U^{\ast}V\otimes I_{H_{A}}$ is an isometry in the multiplier algebra of $C^{\ast}_{f}(X\times Y,A)$. Thus, by Lemma 2.1, $\phi_{\ast}\circ\psi_{\ast}$ is the identity homomorphism on $K_{\ast}(C_{f}^{\ast}(X\times Y,A))$. Similarly, we can define the above two homomorphisms for localization algebras. Therefore, we have the following commutative diagram:
@@
where $N_{X}$ is a locally finite net in $X$. Since the compositions of two vertical homomorphisms are all the identity homomorphism, thus if the middle horizontal map is an isomorphism, then the top horizontal map is injective and the bottom horizontal map is surjective which completes the proof. ∎
@@
Let $X$ be a proper metric space, $\sqcup_{\mathbb{N}}X$ be the separated coarse union of $X$. Then $X$ satisfies the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ if and only if $X\times\square$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$, where $\square$ is the separated coarse union of a sequence of single points.
@@
Firstly, we have that $\sqcup_{\mathbb{N}}X=X\times\square$, thus
@@
Secondly, for an operator $T\in C_{f}^{\ast}(X\times\square,A)$, we have that $\operatorname{prop}_{X}(T)=\operatorname{prop}(T)$, thus
@@
Finally, for a locally finite net $N_{X}$ in $X$, we have that $P_{k}(\sqcup_{\mathbb{N}}X)=\sqcup_{\mathbb{N}}P_{k}(X)$ for any non-negative integer $k$. Thus we have that
$C^{\ast}_{L,f}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C^{\ast}_{L,P_{k}(X),f}(P_{k}(X)\times\square,A);\>\>C_{f}^{\ast}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C_{f}^{\ast}(P_{k}(X)\times\square,A),$
@@
Combing the above lemma with Lemma 4.9, we have the following corollary.
@@
Let $X$ be a proper metric space. If the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $X$, then the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X$.
@@
In [23], the authors introduced the quantitative coarse Baum-Connes conjecture with coefficients which refines the original conjecture and proved this conjecture is equivalent to the uniform coarse Baum-Connes conjecture with coefficients, thus by the above corollary, the coarse Baum-Connes conjecture with filtered coefficients implies that conjecture.
@@ 4.2. Main theorems
A locally finite metric space $N$ is called to have bounded geometry if for every $R>0$, $\sup_{x\in N}\#B(x,R)$ is finite, where #$\#B(x,R)$ is the cardinality of the $R$-ball centered at $x$. A proper metric space $X$ is called to have bounded geometry if there exists a net $N_{X}$ with bounded geometry.
@@
Since $N_{1}$ and $N_{2}$ have bounded geometry, thus the dimension of simplicial complexes $P_{k}(N_{1})$ and $P_{k}(N_{2})$ is less than some positive integer $m_{k}$. Define two maps:
@@
for any $x,y\in P_{k}(N_{1}\times N_{2})$ and $z,w\in P_{k}(N_{1})\times P_{k}(N_{2})$ by direct computations which imply that $\rho$ and $\varphi$ are two coarse and Lipschitz maps. Moreover, $\rho\circ\varphi=id$ and $\varphi\circ\rho$ is strongly Lipschitz homotopic to the identity map by the linear combination of them. Thus, $P_{k}(N_{1}\times N_{2})$ is strongly Lipschitz homotopy equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. And we have that $d(\rho\circ\varphi(z),z)=0$ and $d(\varphi\circ\rho(x),x)\leq 1$, thus $P_{k}(N_{1}\times N_{2})$ is coarsely equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. ∎
@@
$\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{k}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}(P_{k}(N_{X})\times P_{k}(N_{Y}),A)).$
@@
Assume $X$ has bounded geometry. Let $N_{X}$ be a net with bounded geometry in $X$ and $N_{Y}$ be a locally finite net in $Y$. If $Y$ satisfies the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$, then for each $k>0$
@@
$X_{2}=\cup\{\triangle_{2}:\text{$\triangle$ is a simplex of dimension $l$}\},$
@@
    the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $Y$.
@@
Let $N_{X}$, $N_{Y}$ be two nets with bounded geometry in $X$ and $Y$, respectively. Then by Condition 1 and Lemma 3.20, Lemma 4.8, we have that $X\times P_{d}(N_{Y})$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ for any $d\in\mathbb{N}$, namely, the following map is an isomorphism
@@
On the other hand, by Condition 2 and Lemma 4.17, we have the following isomorphism for any $k\in\mathbb{N}$:
@@
$\lim_{k,d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k,d\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times P_{d}(N_{Y}),A)).$
@@
Recall that $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration defined in Lemma 3.9 and $\square$ is the separated coarse union of a sequence of single points. Combining the above theorem with Proposition 4.9 and Corollary 4.12, we have the following theorem which implies that the coarse Baum-Connes conjecture with filtered coefficients is closed under products.
@@
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(Y,A)$ holds for $X$;
@@
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $Y$.
In particular, if the coarse Baum-Connes conjecture with filtered coefficients holds for $X$ and $Y$, then the coarse Baum-Connes conjecture with filtered coefficients holds for $X\times Y$.
When $A=\mathbb{C}$, we have that $C^{\ast}_{f}(\square,\mathbb{C})=\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$ equipped with the trivial filtration, namely, $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))_{r}=\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$ for all $r>0$. Thus, we obtain the following corollary which enlarges the class of product metric spaces that satisfy the coarse Baum-Connes conjecture.
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}(Y)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$.
Then the coarse Baum-Connes conjecture holds for $X\times Y$.

### v2

@@ 4. Product metric spaces
In this section, we will prove that the coarse Baum-Connes conjecture with filtered coefficients for the product metric space of two metric spaces providing each of them satisfies the conjecture with filtered coefficients.
@@ 4.1. The coarse Baum-Connes conjecture along one direction with filtered coefficients
In this subsection, we build a bridge called the localization algebra along one direction with filtered coefficients to connect the coarse Baum-Connes conjecture with filtered coefficients for products and the conjecture with filtered coefficients for their factors.
@@
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two proper metric spaces with two countable dense subsets $Z_{X}$ and $Z_{Y}$, respectively. Let $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. Let $T$ be an operator acting on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$. The propagation along $X$ of $T$, denoted by $\operatorname{prop}_{X}(T)$, is defined to be
@@
Let $X$ and $Y$ be as above, the localization algebra along $X$ with filtered coefficients in $A$ of $X\times Y$, denoted by $C^{\ast}_{L,X,f}(X\times Y,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(X\times Y,A)$ satisfying
@@
Similar definitions as above are also appeared in [2][4].
Let $X$, $X^{\prime}$ and $Y$ be three proper metric spaces. For a uniformly continuous coarse map $g:X\rightarrow X^{\prime}$, then $g\times Id:X\times Y\rightarrow X^{\prime}\times Y$ is also a uniformly continuous coarse map. Thus by Lemma 3.13, there exists a $\ast$-homomorphism $Ad_{g\times Id}:C^{\ast}_{L,f}(X\times Y,A)\rightarrow C^{\ast}_{L,f}(X^{\prime}\times Y,A)$ defined by $V_{g\times Id}=V_{g}\otimes I_{\ell^{2}(Z_{Y})}$. Moreover, for $u\in C^{\ast}_{L,X,f}(X\times Y,A)$, we have that
@@
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2}$ and $X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2}$ and $X_{1}\cap X_{2}$, respectively. Then there is a Mayer-Vietoris six-term exact sequence
${K_{0}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{0}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(C^{\ast}_{L,{X},f}(X\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,{X},f}(X\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{1}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A)),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
Let $X$, $Y$ be two proper metric spaces and $N_{X}$ be a locally finite net in $X$. Then the evaluation at zero map $e$ induces the following isomorphism
@@
Let $g:Y\rightarrow Y^{\prime}$ be a coarsely equivalent map. Then $Id\times g:P_{k}(N_{X})\times Y\rightarrow P_{k}(N_{X})\times Y^{\prime}$ is also coarsely equivalent map, which is uniformly continuous on $P_{k}(N_{X})$. Then by Corollary 3.8, we can show that $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y,A))$ are naturally isomorphic to $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y^{\prime},A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y^{\prime},A))$, respectively, for any $k\in\mathbb{N}$. Thus we completed the proof. ∎
Recall from Lemma 3.9 that $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with the filtration induced by the propagation on $Y$ and the filtration on $A$. Then we have the following relation between the coarse Baum-Connes conjecture along $X$ with filtered coefficients and the coarse Baum-Connes conjecture with filtered coefficients.
@@
Let $H$ be a separable Hilbert space and $U:H\rightarrow H\otimes H$ be a unitary. Fixed a unit vector $h_{0}\in H$ and let $P:H\rightarrow H,h\mapsto\langle h,h_{0}\rangle h_{0}$. Then $P$ is a rank-one projection. Recall that the algebra $C_{f}^{\ast}(X\times Y,A)$ acts on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ and the algebra $C^{\ast}_{f}(X,C^{\ast}(Y,A))$ acts on $\ell^{2}(Z_{X})\otimes H\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$, where $Z_{X}$ and $Z_{Y}$ are two countable dense subsets in $X$ and $Y$, respectively. Define two $\ast$-homomorphisms:
@@
where $V:H\rightarrow H\otimes H$ maps $h$ to $h\otimes h_{0}$. Since the operator $I_{\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})}\otimes U^{\ast}V\otimes I_{H_{A}}$ is an isometry in the multiplier algebra of $C^{\ast}_{f}(X\times Y,A)$. Thus, by Lemma 2.1, $\phi_{\ast}\circ\psi_{\ast}$ is the identity homomorphism on $K_{\ast}(C_{f}^{\ast}(X\times Y,A))$. Moreover, we can similarly define two homomorphisms between $C^{\ast}_{L,f}(X,C^{\ast}(Y,A))$ and $C_{L,X,f}^{\ast}(X\times Y,A)$, and obtain the similar result as above. Therefore, we have the following commutative diagram:
@@
where $N_{X}$ is a locally finite net in $X$. Since the compositions of two vertical homomorphisms are all the identity homomorphism. Thus if the middle horizontal map is an isomorphism, then the top horizontal map is injective and the bottom horizontal map is surjective. This implies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ holds for $X\times Y$. ∎
@@
Finally, we simplify the consistent coarse Baum-Connes conjecture with filtered coefficients in this subsection.
@@
Let $X$ be a proper metric space, $\sqcup_{\mathbb{N}}X$ be the separated coarse union of $X$. Then $X$ satisfies the consistent coarse Baum-Connes conjecture with filtered coefficients in $A$ if and only if $X\times\square$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$, where $\square$ is the separated coarse union of a sequence of single points.
@@
Firstly, we have that $\sqcup_{\mathbb{N}}X=X\times\square$. Thus
@@
Secondly, for an operator $T\in C_{f}^{\ast}(X\times\square,A)$, we have that $\operatorname{prop}_{X}(T)=\operatorname{prop}(T)$. Thus
@@
Finally, for a locally finite net $N_{X}$ in $X$, we have that $P_{k}(\sqcup_{\mathbb{N}}X)=\sqcup_{\mathbb{N}}P_{k}(X)$ for any non-negative integer $k$. Thus we have
$\begin{split}&C^{\ast}_{L,f}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C^{\ast}_{L,P_{k}(X),f}(P_{k}(X)\times\square,A);\\ &C_{f}^{\ast}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C_{f}^{\ast}(P_{k}(X)\times\square,A),\end{split}$
@@
Definition 4.12.
[17, Definition 5.14] Let $\{A^{i}\}_{i\in\mathcal{I}}$ be a family of filtered $C^{\ast}$-algebra. The uniform product of $\{A^{i}\}_{i\in\mathcal{I}}$, denoted by $\prod^{u}_{i\in\mathcal{I}}A^{i}$, is defined to be the norm closure of the set $\bigcup_{r>0}(\prod_{i\in\mathcal{I}}A^{i}_{r})$ in the direct product $\prod_{i\in\mathcal{I}}A^{i}$.
Remark 4.13.
Let $\Gamma$ be a finitely generated group equipped with the word length metric. Then the Roe algebra $C^{\ast}(\Gamma)$ is $\ast$-isomorphic to the reduced crossed product $\ell^{\infty}(\Gamma,\mathfrak{K}(H))\rtimes_{r}\Gamma$ (please see [1, Proposition 5.1.3] for the proof). By a similar argument, the Roe algebra $C^{\ast}_{f}(\Gamma,A)$ with filtered coefficients in $A$ is $\ast$-isomorphic to the reduced crossed product $(\prod^{u}_{\Gamma}(A\otimes\mathfrak{K}(H)))\rtimes_{r}\Gamma$, where $A\otimes\mathfrak{K}(H)$ is a filtered $C^{\ast}$-algebra equipped with the filtration $(A_{r}\otimes\mathfrak{K}(H))_{r>0}$.
Let $A$ be a filtered $C^{\ast}$-algebra. Then, the uniform product $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))$ is a filtered $C^{\ast}$-algebra equipped with the filtration $(\prod_{\mathbb{N}}(A_{r}\otimes\mathfrak{K}(H)))_{r>0}$. And we have that
$C^{\ast}_{f}(\square,A)=\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H)),$
where $\square$ is the separated coarse union of a sequence of single points.
Thus, combing Lemma 4.11 with Proposition 4.9, we obtain the following corollary.
@@
Let $X$ be a proper metric space. If the coarse Baum-Connes conjecture with filtered coefficients in $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))$ holds for $X$, then the consistent coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X$.
@@
In [26], the authors introduced the quantitative coarse Baum-Connes conjecture with coefficients and proved that this conjecture is equivalent to the consistent coarse Baum-Connes conjecture with coefficients. Thus by the above corollary, the coarse Baum-Connes conjecture with filtered coefficients implies that conjecture.
@@ 4.2. Main theorems
A locally finite metric space $N$ is called to have bounded geometry, if $\sup_{x\in N}\#B(x,R)$ is finite for each $R>0$, where #$\#B(x,R)$ is the cardinality of the $R$-ball centered at $x$. A proper metric space $X$ is called to have bounded geometry, if there exists a net $N_{X}$ with bounded geometry.
@@
Since $N_{1}$ and $N_{2}$ have bounded geometry, thus the dimensions of simplicial complexes $P_{k}(N_{1})$ and $P_{k}(N_{2})$ are less than some positive integer $m_{k}$. Define two maps:
@@
for any $x,y\in P_{k}(N_{1}\times N_{2})$ and $z,w\in P_{k}(N_{1})\times P_{k}(N_{2})$ by direct computations, which imply that $\rho$ and $\varphi$ are two coarse and Lipschitz maps. Moreover, $\rho\circ\varphi=id$ and $\varphi\circ\rho$ is strongly Lipschitz homotopic to the identity map by the linear combination of them. Thus, $P_{k}(N_{1}\times N_{2})$ is strongly Lipschitz homotopy equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. And we have that $d(\rho\circ\varphi(z),z)=0$ and $d(\varphi\circ\rho(x),x)\leq 1$, thus $P_{k}(N_{1}\times N_{2})$ is coarsely equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. ∎
@@
$\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{k}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times P_{k}(N_{Y}),A)).$
@@
Assume $X$ has bounded geometry. Let $N_{X}$ be a net with bounded geometry in $X$ and $N_{Y}$ be a locally finite net in $Y$. If $Y$ satisfies the consistent coarse Baum-Connes conjecture with filtered coefficients in $A$, then for each $k>0$
@@
$X_{2}=\cup\{\triangle_{2}:\text{$\triangle$ is a simplex of dimension $l$}\}.$
@@
    the consistent coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $Y$.
@@
Let $N_{X}$, $N_{Y}$ be two nets with bounded geometry in $X$ and $Y$, respectively. Then combining Condition (1) with Lemma 3.20 and Lemma 4.8, we have that $X\times P_{d}(N_{Y})$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ for any $d\in\mathbb{N}$, namely, the following map is an isomorphism
@@
On the other hand, by Condition (2) and Lemma 4.19, we have the following isomorphism for any $k\in\mathbb{N}$:
@@
$\lim_{k,d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k,d\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times P_{d}(N_{Y}),A)),$
@@
Let $A$ be a filtered $C^{\ast}$-algebra. Recall that the Roe algebra $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with the filtration $(C^{\ast}_{f}(Y,A))_{r>0}$ defined in Lemma 3.9 and the uniform product $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))$ is also a filtered $C^{\ast}$-algebra equipped with the filtration $(\prod_{\mathbb{N}}(A_{r}\otimes\mathfrak{K}(H)))_{r>0}$. Combining Theorem 4.20 with Proposition 4.9 and Corollary 4.14, we have the following theorem which implies that the coarse Baum-Connes conjecture with filtered coefficients is closed under products.
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the Roe algebra $C^{\ast}_{f}(Y,A)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the uniform product $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))$.
Then $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$.
In particular, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture with filtered coefficients, then $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients.
Remark 4.22.
For two discrete groups $\Gamma_{1}$ and $\Gamma_{2}$, Oyono-Oyono proved that $\Gamma_{1}\times\Gamma_{2}$ satisfies the Baum-Connes conjecture with coefficients if and only if $\Gamma_{1}$ and $\Gamma_{2}$ satisfy the Baum-Connes conjecture with coefficients in [15, Corollary 7.12]. Thus, we can see Theorem 4.21 as a coarse analogue of Oyono-Oyono’s result. Besides, for any proper metric space $X$, the product $X\times[0,\infty)$ always satisfies the coarse Baum-Connes conjecture with filtered coefficients by an argument of the Eilenberg swindle (see [21, Proposition 7.5.2]). However, we know that the coarse Baum-Connes conjecture with filtered coefficients does not hold for some expanders. Thus, Theorem 4.21 is not a necessary and sufficient condition in general.
When $A=\mathbb{C}$, we have that $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))=\prod_{\mathbb{N}}\mathfrak{K}(H)$ equipped with the trivial filtration, namely, $(\prod_{\mathbb{N}}\mathfrak{K}(H))_{r}=\prod_{\mathbb{N}}\mathfrak{K}(H)$ for each $r>0$. Thus, we obtain the following corollary which enlarges the class of product metric spaces that satisfy the coarse Baum-Connes conjecture.
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the Roe algebra $C^{\ast}(Y)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in $\prod_{\mathbb{N}}\mathfrak{K}(H)$.
Then $X\times Y$ satisfies the coarse Baum-Connes conjecture.
@@
Remark 4.24.
For any $p\in[1,\infty)$, there is a $\ell^{p}$ coarse Baum-Connes conjecture for computing the $K$-theory of $\ell^{p}$ Roe algebras (see [28]). Actually, Theorem 4.21 and Corollary 4.23 also hold for the $\ell^{p}$ coarse Baum-Connes conjecture by the similar proofs.